OpenStudy (anonymous):

what is the point slope equation of the line with slope 3/2 that goes thru the point (-3,5)

OpenStudy (anonymous):

y=mx+c

OpenStudy (anonymous):

y=0.5x+6.5

OpenStudy (karatechopper):

@Raja99 , what you wrote it slope intercept form. Point-slope form is: y-y1=m(x-x1)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

so...its y - 5 = -3/2(x+3) ?

OpenStudy (anonymous):

ya

OpenStudy (jennychan12):

use point slope form. easier Y - Y1 = m(X-X1) Y - 5 = 3/2(X+3) solve for y. Y -5 = 3/2x-9/2 Y = 3/2+34/5